# Bridging constrained random-phase approximation and linear response theory for computing Hubbard parameters

### v1

@@ Abstract
The very small values for $U_{\text{cRPA}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$ obtained for the $d$-only case are related to the strong hybridization between the interacting and screening subspaces in the underlying bands. Several methods have been suggested obtain $\chi_{0}^{\mathcal{R}}$ in such cases [48, 65, 66]. Here, we use the method introduced in Ref. [65], which assigns weights to each particle-hole transition in Eq. (2) according to the product of the weights of the $\mathcal{D}$ subspace in the corresponding Bloch functions. For bands with mixed character that are close to or even crossing the Fermi level, this can lead to transitions with very low energy which contribute strongly to the screening and thus lead to a very small $U_{\text{cRPA}}$ [^2] We note that even though the “projector method” [66] (not implemented in RESPACK) typically results in larger interaction parameters compared to the “weighted method” (see, e.g., [59, 67]), it nevertheless suffers from the same problem of very low energy transitions and a resulting overscreening [59]. This behavior is closely related to the general tendency of cRPA to overscreen [68, 69, 70, 71], which is strongest if the interacting and screening subspaces are close in energy, such that corrections beyond RPA become important [68]. The $d$-only case thus represents a problematic case, where the general applicability of the cRPA approach becomes questionable, and the results can furthermore depend sensitively on the specific way the two subspaces are separated. On the other hand, the $d$-only case is quite relevant for practical applications, since in DFT+$U$ and DFT+DMFT calculations for TM compounds, one typically applies the $U$ correction only on the TM sites, while the ligand states are considered as weakly interacting. Interestingly, it was recently noted that explicitly considering interaction effects also on the ligands can in fact improve the quantitative description of various TM compounds [72].
@@
Within LRT, apart from the fact that the $d$-only MLWFs are slightly more localized, the main difference to the $d$-$p$ case is the absence of TM–ligand off-diagonal elements in the susceptibility matrices $\tilde{\chi}^{IJ}$ and $\tilde{\chi}_{0}^{IJ}$, which affect $U^{I}_{\text{LRT}}$ due to the inversion in Eq. (1). These off-diagonal elements are larger in Sr2FeO4 than in KCuF3, due to the stronger $d$-$p$ hybridization in the oxide compared to the more ionic fluoride, which results in a larger difference of $U_{\text{LRT}}$ between the two localized basis sets for Sr2FeO4. However, even though the very strong overscreening observed in cRPA for the $d$-only case seems to be absent in LRT, we note that, since the generalized susceptibility $(\chi_{0})_{ijkl}^{IJKL}$ is equivalent to the bare cRPA susceptibility [64], the general overscreening effects identified for cRPA [68, 69, 70, 71] can be expected to also affect $U_{\text{LRT}}$, even though the practical problem of how to separate two strongly hybridized subspaces does not appear. This could explain why, in the case of Sr2FeO4, for which the frontier bands overlap slightly with the lower-lying screening bands, both $U_{\text{LRT}}$ and $U_{\text{cRPA}}-J_{\text{cRPA}}$ in the frontier basis are too low to reproduce the experimentally observed insulating state [42].
In summary, we have incorporated the use of Wannier functions as Hubbard projectors in LRT calculation of the Hubbard $U$, enabling a direct quantitative comparison with corresponding cRPA calculations. Our numerical results show that in cases where the interacting subspace corresponds to an isolated set of bands, both LRT and cRPA yield comparable but not identical Hubbard parameters. We further showed, that the quantitative agreement can be understood also on a formal level, since both methods are based on similar approximations. Differences arise from neglecting the change of the exchange-correlation potential in the cRPA response and the fact that LRT takes into account additional screening effects related to local exchange as well as inter-site and inter-orbital contributions within the interacting subspace.
For cases where the interacting and screening subspaces are strongly hybridized and do not correspond to a separated set of bands, we find that cRPA can produce unphysically small interaction values, questioning its applicability in such cases. While $U_{\text{LRT}}$ appears unaffected by this problem, the fact that LRT does not systematically exclude screening within the interacting subspace could lead to a “double-counting” in the subsequent treatment of interaction effects within the interacting subspace.
@@
We are thankful to Matteo Cococcioni and Kazuma Nakamura for valuable discussions. This research was supported by ETH Zürich. The Flatiron Institute is a division of the Simons Foundation. I.T. acknowledges partial support by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (Grant number 205602). Calculations were performed on the “Euler” cluster of ETH Zürich.

### v2

@@ Abstract
We now turn to the localized $d$-only case (last set of bars in both panels of Fig. 2), for which the interacting and screening subspaces are strongly hybridized. Several methods have been suggested to obtain $\chi_{0}^{\mathcal{R}}$ in such cases [miyake_initio_2009, sasioglu_effective_2011, Kaltak:2015, Kaltak_spectral_2025]. Here, we use the weighted method introduced in Ref. [sasioglu_effective_2011]. While $U_{\text{cRPA}}$ is very small in this case, $U_{\text{LRT}}$ is much larger and comparable to the values obtained for the localized $d$-$p$ basis. The very small $U_{\text{cRPA}}$ for the $d$-only case is due to the strong screening from transitions with very low energy, related to bands with mixed character that are close to or even crossing the Fermi level. The $d$-only case thus represents a notoriously problematic case, where the general applicability of the cRPA approach becomes questionable, andthe results can depend sensitively on the specific way the two subspaces are separated.  [^2] We note that even though the “projector method” [Kaltak:2015] (not implemented in RESPACK) typically results in larger interaction parameters compared to the weighted method (see, e.g., [merkel_calculation_2024a, Reddy/Kaltak/Kim:2025]), it nevertheless suffers from the same problem [merkel_calculation_2024a]. On the other hand, the $d$-only case is quite relevant for practical applications, since in DFT+$U$ and DFT+DMFT calculations for TM compounds, one typically applies the $U$ correction only on the TM sites, while the ligand states are considered as weakly interacting.
@@
Within LRT, apart from the fact that the $d$-only MLWFs are slightly more localized, the main difference to the $d$-$p$ case is the absence of TM–ligand off-diagonal elements in the susceptibility matrices $\tilde{\chi}^{IJ}$ and $\tilde{\chi}_{0}^{IJ}$, which affect $U^{I}_{\text{LRT}}$ due to the inversion in Eq. (4). These off-diagonal elements are larger in Sr2FeO4 than in KCuF3, due to the stronger $d$-$p$ hybridization in the oxide compared to the more ionic fluoride, which results in a larger difference of $U_{\text{LRT}}$ between the two localized basis sets for Sr2FeO4.
We note that, since the generalized susceptibility $(\chi_{0})_{ijkl}^{IJKL}$ is equivalent to the bare cRPA susceptibility [SI] (for clearly separable $\mathcal{D}$), the well known overscreening effects identified for cRPA when interacting and screening subspaces are close in energy or overlap [vanloon_random_2021, honerkamp_limitations_2018a, werner_dynamical_2016, shinaoka_accuracy_2015, Chang2024], can be expected to also affect $U_{\text{LRT}}$ in such cases. As noted recently, such overscreening can also be due to an incorrect energy separation between the frontier and ligand states in the underlying band structure  [Carta/Panda/Ederer:2025]. This could explain why in the Sr2FeO4 frontier basis, for which $\mathcal{D}$ overlaps slightly with the lower-lying screening bands, $U_{\text{LRT}}$ and $U_{\text{cRPA}}$ are too low to reproduce the experimentally observed insulating state [kazemi-moridani_strontium_2024a].
In summary, we have incorporated the use of Wannier functions as Hubbard projectors in LRT calculation of the Hubbard $U$, enabling a direct quantitative comparison with corresponding cRPA calculations. Our analysis shows that, in cases where the interacting subspace corresponds to an isolated set of bands, both approaches can be understood within the same framework. Differences in the corresponding $U$ values are related to neglecting the xc contribution to the response within cRPA and the coarse-graining to purely monopolar response functions within LRT, which effectively incorporates additional excitation channels within the interacting subspace. Our numerical results show that, if these differences are appropriately taken into account, excellent quantitative agreement can be achieved, thus solving the long standing question about the equivalence of the two methods. For cases where the interacting and screening subspaces do not correspond to a separated set of bands, we find that cRPA produces unphysically small interaction values, questioning its applicability in such cases.
In practice, given the close relation between the two methods, we believe that choosing one over the other depends mostly on what kind of calculation (DFT+$U$, DFT+DMFT, etc.) is of interest and whether an explicit treatment of the effects coming from the xc contribution or the exclusion of the multipolar screening channels is considered to be important.
@@
We are thankful to Matteo Cococcioni and Kazuma Nakamura for valuable discussions. This research was supported by ETH Zürich. The Flatiron Institute is a division of the Simons Foundation. I.T. acknowledges partial support by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (SNSF) Grant No. 205602, and SNSF Grants No. 200021-227641 and No. 200021-236507. Calculations were performed on the “Euler” cluster of ETH Zürich.